# Networks of neural networks: more is different

### v1

@@ Appendix A RS solution by interpolation technique
Exploiting once more a $L-$dimensional multivariate Gaussian transform, we can linearize the last term of the argument of the exponential function in (A.14) and explicitly perform the sum over $\{\sigma^{(a)}\}$, getting the one-body statistical pressure
$\begin{array}[]{lll}\mathcal{A}(\beta,H,\bm{g},\bm{h};t=0)&=&-\dfrac{\beta\gamma}{2}+L\log{2}+\displaystyle\sum\limits_{a=1}^{L}\mathbb{E}_{\bm{\xi},x}\cosh\Bigg{(}\left[\displaystyle\sum\limits_{\mu=1}^{L}\beta\xi^{\mu}\displaystyle\sum\limits_{b=1}^{L}g_{ab}\bar{m}_{\mu}^{b}+\beta Hh^{a}(t)+x\sqrt{\beta\gamma\bar{p}_{12}^{a}}\right]\Bigg{)}\\ \\ &&-\dfrac{1}{2}\log\Big{[}\mathrm{det}\mathcal{V}\Big{]}-\dfrac{\gamma}{2}\log\Big{[}\mathrm{det}\,\mathcal{G}\Big{]}+\dfrac{\beta\gamma}{2}\displaystyle\sum\limits_{a,b=1}^{L}\sqrt{\bar{q}_{12}^{a}}(\mathcal{G}^{-1}\,)_{ab}\,\sqrt{\bar{q}_{12}^{b}}\end{array}$ (A.16)
@@
$\begin{array}[]{lll}\displaystyle\int\mathcal{D}(\tau_{a})=\displaystyle\int\prod\limits_{b=1}^{L}\dfrac{d\tau_{a}d\tau_{b}}{2\pi}\exp\left(-\dfrac{1}{2}\displaystyle\sum\limits_{b=1}^{L}\tau_{a}(\mathcal{V}^{-1})_{ab}\,\tau_{b}\right)\end{array}$ (A.17)
@@
$\begin{array}[]{lll}\mathcal{A}(\beta,H,\bm{g},\bm{h})&=&-\dfrac{\beta\gamma}{2}+L\log{2}+\displaystyle\sum\limits_{a=1}^{L}\mathbb{E}_{\bm{\xi},x}\cosh\Bigg{(}\left[\displaystyle\sum\limits_{\mu=1}^{L}\beta\xi^{\mu}\displaystyle\sum\limits_{b=1}^{L}g_{ab}\bar{m}_{\mu}^{b}+\beta Hh^{a}(t)+x\sqrt{\beta\gamma\bar{p}_{12}^{a}}\right]\Bigg{)}\\ \\ &&-\dfrac{1}{2}\log\Big{[}\mathrm{det}\mathcal{V}\Big{]}-\dfrac{\gamma}{2}\log\Big{[}\mathrm{det}\,\mathcal{G}\Big{]}+\dfrac{\beta\gamma}{2}\displaystyle\sum\limits_{a,b=1}^{L}\sqrt{\bar{q}_{12}^{a}}(\mathcal{G}^{-1}\,)_{ab}\,\sqrt{\bar{q}_{12}^{b}}\\ \\ &&-\dfrac{\beta}{2}\displaystyle\sum\limits_{a,b=1}^{L}\displaystyle\sum\limits_{\mu=1}^{L}\bar{m}_{\mu}^{a}g_{ab}\bar{m}_{\mu}^{b}-\dfrac{\beta\gamma}{2}\displaystyle\sum\limits_{a\neq b}\Big{(}\bar{p}_{11}^{ab}\bar{q}_{11}^{ab}-\bar{p}_{12}^{ab}\bar{q}_{12}^{ab}\Big{)}-\dfrac{\beta\gamma}{2}\displaystyle\sum\limits_{a}\bar{p}_{12}^{a}\Big{(}1-\bar{q}_{12}^{a}\Big{)}.\end{array}$ (A.18)
The previous expression can be further simplified by noting that its extremization with respect to $\bar{q}_{11}^{ab}$ and $\bar{q}_{12}^{ab}$ yields the following relations:
@@
$\begin{array}[]{lll}\mathcal{A}(\beta,H,\bm{g},\bm{h})&=&L\log 2+\displaystyle\sum\limits_{a=1}^{L}\mathbb{E}_{\bm{\xi},x}\log\Bigg{\{}\cosh\Bigg{[}\displaystyle\sum\limits_{\mu=1}^{L}\beta\xi^{\mu}\displaystyle\sum\limits_{b=1}^{L}g_{ab}\bar{m}_{\mu}^{b}+\beta Hh^{a}(t)+x\sqrt{\beta\gamma\bar{p}_{12}^{a}}\Bigg{]}\Bigg{\}}\\ \\ &&-\dfrac{\gamma}{2}\log\Big{[}\mathrm{det}\mathcal{G}\Big{]}+\dfrac{\beta\gamma}{2}\displaystyle\sum\limits_{a,b=1}^{L}\sqrt{\bar{q}_{12}^{a}}(\mathcal{G}^{-1}\,)_{ab}\,\sqrt{\bar{q}_{12}^{b}}\\ \\ &&-\dfrac{\beta}{2}\displaystyle\sum\limits_{a,b=1}^{L}\displaystyle\sum\limits_{\mu=1}^{L}\bar{m}_{\mu}^{a}g_{ab}\bar{m}_{\mu}^{b}-\dfrac{\beta\gamma}{2}\displaystyle\sum\limits_{a}\bar{p}_{12}^{a}\Big{(}1-\bar{q}_{12}^{a}\Big{)}\end{array}$ (A.20)
@@
$\begin{array}[]{lll}\bar{p}_{12}^{c}&=&\dfrac{1}{\beta}\dfrac{\partial_{\bar{q}_{12}^{c}}\Big{[}\mathrm{det}\mathcal{G}\Big{]}}{\mathrm{det}\mathcal{G}}-\partial_{\bar{q}_{12}^{c}}\left[\displaystyle\sum\limits_{a,b=1}^{L}\sqrt{\bar{q}_{12}^{a}}\;(\mathcal{G}^{-1}\,)_{ab}\,\sqrt{\bar{q}_{12}^{b}}\right].\end{array}$ (A.24)
@@ Appendix B Low-load self-consistency equations for $L=3$
$\begin{array}[]{lll}\mathcal{M}=\begin{pmatrix}1-(1-\lambda^{2})\displaystyle\sum\limits_{i\neq 1}^{3}d^{i}+\mathrm{det}g\prod\limits_{i\neq 1}^{3}d^{i}&-\lambda[1-(1+\lambda)d^{3}]&-\lambda[1-(1+\lambda)d^{2}]\\ -\lambda[1-(1+\lambda)d^{3}]&1-(1-\lambda^{2})\displaystyle\sum\limits_{i\neq 2}^{3}d^{i}+\mathrm{det}g\prod\limits_{i\neq 2}^{3}d^{i}&-\lambda[1-(1+\lambda)d^{1}]\\ -\lambda[1-(1+\lambda)d^{2}]&-\lambda[1-(1+\lambda)d^{1}]&1-(1-\lambda^{2})\displaystyle\sum\limits_{i\neq 3}^{3}d^{i}+\mathrm{det}g\prod\limits_{i\neq 3}^{3}d^{i}\end{pmatrix}.\end{array}$ (B.6)
@@
Figure 10: The region in the plane $(\beta,\lambda)$ where the five-layer model is expected to successfully disentangle mixtures of five patterns is depicted by solving the self-consistent equations (11) (dashed lines) and compared to MC simulations (color map), for several values of the external field: $H=0.0$ (left column), $H=0.1$ (middle column) and $H=0.2$ (right column), and two different thresholds on the magnetizations $m>0.95$ (upper row), $m>0.99$ (lower row), in analogy to Fig. 5. The self-consistency equations have been solved in the $\gamma=0$ case, while the disentangling accuracy has been computed by averaging over $50$ statistically-independent MC runs, each with $N=5000$ and $K=5$. In each run the model is initialized in the $\bm{\sigma}^{(h)}$ configuration and let evolve up to equilibrium; the final magnetizations have been obtained by computing the overlap between the state of each layer and the five patterns $\bm{\xi}^{1},..,\bm{\xi}^{5}$.

### v2

@@ Appendix A RS solution by interpolation technique
Exploiting once more a $L$-dimensional multivariate Gaussian transform, we can linearize the last term of the argument of the exponential function in (A.14) and explicitly perform the sum over $\{\sigma^{a}\}$, getting the one-body statistical pressure
$\begin{array}[]{lll}\mathcal{A}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}RS}}(\beta,H,\bm{g},\bm{h};t=0)&=&-\dfrac{\beta\gamma}{2}+L\log{2}+\displaystyle\sum\limits_{a=1}^{L}\mathbb{E}_{\bm{\xi},x}\cosh\Bigg(\left[\displaystyle\sum\limits_{\mu=1}^{L}\beta\xi^{\mu}\displaystyle\sum\limits_{b=1}^{L}g_{ab}\bar{m}_{\mu}^{b}+\beta Hh^{a}(t)+x\sqrt{\beta\gamma\bar{p}_{12}^{a}}\right]\Bigg)\\ \\ &&-\dfrac{1}{2}\log\Big[\mathrm{det}\mathcal{V}\Big]-\dfrac{\gamma}{2}\log\Big[\mathrm{det}\,\mathcal{G}\Big]+\dfrac{\beta\gamma}{2}\displaystyle\sum\limits_{a,b=1}^{L}\sqrt{\bar{q}_{12}^{a}}(\mathcal{G}^{-1}\,)_{ab}\,\sqrt{\bar{q}_{12}^{b}}\end{array}$ (A.16)
@@
$\begin{array}[]{lll}\displaystyle\int\mathcal{D}(\tau)=\displaystyle\int\prod\limits_{b=1}^{L}\dfrac{d\tau_{a}d\tau_{b}}{2\pi}\exp\left(-\dfrac{1}{2}\displaystyle\sum\limits_{b=1}^{L}\tau_{a}(\mathcal{V}^{-1})_{ab}\,\tau_{b}\right)\end{array}$ (A.17)
@@
$\begin{array}[]{lll}\mathcal{A}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}RS}}(\beta,H,\bm{g},\bm{h})&=&-\dfrac{\beta\gamma}{2}+L\log{2}+\displaystyle\sum\limits_{a=1}^{L}\mathbb{E}_{\bm{\xi},x}\cosh\Bigg(\left[\displaystyle\sum\limits_{\mu=1}^{L}\beta\xi^{\mu}\displaystyle\sum\limits_{b=1}^{L}g_{ab}\bar{m}_{\mu}^{b}+\beta Hh^{a}(t)+x\sqrt{\beta\gamma\bar{p}_{12}^{a}}\right]\Bigg)\\ \\ &&-\dfrac{1}{2}\log\Big[\mathrm{det}\mathcal{V}\Big]-\dfrac{\gamma}{2}\log\Big[\mathrm{det}\,\mathcal{G}\Big]+\dfrac{\beta\gamma}{2}\displaystyle\sum\limits_{a,b=1}^{L}\sqrt{\bar{q}_{12}^{a}}(\mathcal{G}^{-1}\,)_{ab}\,\sqrt{\bar{q}_{12}^{b}}\\ \\ &&-\dfrac{\beta}{2}\displaystyle\sum\limits_{a,b=1}^{L}\displaystyle\sum\limits_{\mu=1}^{L}\bar{m}_{\mu}^{a}g_{ab}\bar{m}_{\mu}^{b}-\dfrac{\beta\gamma}{2}\displaystyle\sum\limits_{\begin{subarray}{c}a,b=1\\ a\neq b\end{subarray}}^{L}\Big(\bar{p}_{11}^{ab}\bar{q}_{11}^{ab}-\bar{p}_{12}^{ab}\bar{q}_{12}^{ab}\Big)-\dfrac{\beta\gamma}{2}\displaystyle\sum\limits_{a=1}^{L}\bar{p}_{12}^{a}\Big(1-\bar{q}_{12}^{a}\Big).\end{array}$ (A.18)
Since stationary configurations correspond to those values of the order parameters that maximize the statistical pressure (or equivalently, minimize the free energy) of the system, and in this analysis we are only interested in the values of the order parameters that are saddle points of the free energy (A.18), the previous expression can be further simplified by observing that its extremization with respect to $\bar{q}_{11}^{ab}$ and $\bar{q}_{12}^{ab}$ leads to the following relations:
@@
$\begin{array}[]{lll}\mathcal{A}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}RS}}(\beta,H,\bm{g},\bm{h})&=&L\log 2+\displaystyle\sum\limits_{a=1}^{L}\mathbb{E}_{\bm{\xi},x}\log\Bigg\{\cosh\Bigg[\displaystyle\sum\limits_{\mu=1}^{L}\beta\xi^{\mu}\displaystyle\sum\limits_{b=1}^{L}g_{ab}\bar{m}_{\mu}^{b}+\beta Hh^{a}(t)+x\sqrt{\beta\gamma\bar{p}_{12}^{a}}\Bigg]\Bigg\}\\ \\ &&-\dfrac{\gamma}{2}\log\Big[\mathrm{det}\mathcal{G}\Big]+\dfrac{\beta\gamma}{2}\displaystyle\sum\limits_{a,b=1}^{L}\sqrt{\bar{q}_{12}^{a}}(\mathcal{G}^{-1}\,)_{ab}\,\sqrt{\bar{q}_{12}^{b}}\\ \\ &&-\dfrac{\beta}{2}\displaystyle\sum\limits_{a,b=1}^{L}\displaystyle\sum\limits_{\mu=1}^{L}\bar{m}_{\mu}^{a}g_{ab}\bar{m}_{\mu}^{b}-\dfrac{\beta\gamma}{2}\displaystyle\sum\limits_{a=1}^{L}\bar{p}_{12}^{a}\Big(1-\bar{q}_{12}^{a}\Big)\end{array}$ (A.20)
@@
$\begin{array}[]{lll}\bar{p}_{12}^{c}&=&\dfrac{1}{\beta}\dfrac{\partial_{\bar{q}_{12}^{c}}\Big[\mathrm{det}\mathcal{G}\Big]}{\mathrm{det}\mathcal{G}}-\partial_{\bar{q}_{12}^{c}}\left[\displaystyle\sum\limits_{a,b=1}^{L}\sqrt{\bar{q}_{12}^{a}}\;(\mathcal{G}^{-1}\,)_{ab}\,\sqrt{\bar{q}_{12}^{b}}\right],\end{array}$ (A.24)
which are obtained by the extremization of the (A.20) with respect to $\bar{m}_{\nu}^{a},\bar{q}_{12}^{a}$ and $\bar{p}_{12}^{c}$.
In this work we will specialize on the low-load regime, i.e. $\gamma=0$, where the RS assumption is exact, much as like the standard Hopfield model, e.g., see Bovier . On the other hand, in the high-load regime $\gamma\in\mathbb{R}^{+}$, Replica-Symmetry-Breaking (RSB) phenomena are expected to emerge and their onset can, for instance, be addressed by determining the so-called de Almeida-Thouless line, e.g., see AT-Linda , that traces -in the space of the control parameters- the boundaries of the stability of the RS solution.
@@ Appendix B Low-load self-consistency equations for $L=3$
[…]